solve the equation 9=2^x by applying the natural logarithm function, log.e to both sides
is it a case of x=log9/log2 ???

Answer by Alan3354(69443)

solve the equation 9=2^x by applying the natural logarithm function, log.e to both sides
is it a case of x=log9/log2 ???
---------------
That's it, except for using ln
-----
x = ln(9)/ln(2)
The answer is the same, x =~ 3.1699